OpenStudy (anonymous):

find the vertex focus and directrix of the parabola x = 1/4 (y^2 + 2y + 33)

OpenStudy (anonymous):

The standard form of the parabola equation is \[(y-k)^{2}=4p(x-h)\] so transforming x = 1/4 (y^2 + 2y + 33) to: x=1/4[(y^2 + 2y + 1) + 32)] x=14[ (y+1)^2 + 32] \[(y+1)^2 = 4*1*(x-8)\] so k = -1. h = 8 , p =1 1. vertex is (h,k) = (8,-1) 2. focus is (h,k+p) = (8,-1+1) = (8,0) 3. directrix is the equation y=k-p, so y=-2